# Distance-Domain Degrees of Freedom in Near-Field Region

### v3

@@ Abstract
Extremely large aperture array operating in the near‐field regime unlock additional spatial resources, which can be exploited to simultaneously serve multiple users even when they share the same angular direction. This work investigate the distance-domain degrees of freedom (DoF), which are the DoF when the user only varies its distance to the base station and not the angle. To obtain the distance-domain DoF, we investigate a line-of-sight (LoS) channel of a source representing a base station and an observation region representing users, where the source is a large two-dimensional transmit (Tx) array with arbitrary shape and the observation region is an arbitrarily long linear receive (Rx) array with collinearly-aligned elements but located at different distances from the Tx array. We assume that both the Tx and Rx array has continuous aperture with an infinite number of elements and infinitesimal spacing, which establishes an upper bound for the distance-domain DoF in the case of finite number of elements. First, we assume an ideal case where the Tx array is a single piece and the Rx array is on the broad side of the Tx array. By reformulating the channel as an integral operator with a Hermitian convolution kernel, we derive a closed-form expression for the distance-domain DoF via the Fourier transform. Our analysis shows that the distance-domain DoF is predominantly determined by the extreme boundaries of both the Tx and Rx arrays rather than its detailed interior structure. We further extend the framework to non-broadside configurations by employing a projection method that converts the problem to an equivalent broadside case. Finally, we extend our analytical framework to the modular array, which shows the distance-domain DoF gain over the single-piece array given the constraint of the physical length of the array.
@@ I Introduction
Unlike far-field scenarios, where beamforming is mainly used to steer energy along specific angular directions, the near-field enables beam focusing that leverages both angular and distance domains. This capability permits spatial multiplexing of users that share the same angular direction if they are located at different distances. Such an approach could be particularly useful in ultra-dense scenarios where many users are aligned in the same direction, a situation that challenges traditional far-field beamforming methods. Several studies have investigated spatial multiplexing based on the distance domain; for instance, [zhang2022beam] and [li2023near] demonstrated through simulations that two collinear users can be spatially multiplexed in a LoS near-field channel. However, these works did not delve into a theoretical explanation of channel characteristics for the effectiveness of such multiplexing. To address this gap, [wu2023multiple] and [wu2023enabling] showed that the channel correlation between collinear users tends to decrease as their separation increases in the near-field region, so that once the correlation falls below a certain threshold, multiplexing becomes feasible. Based on the closed‐form channel‐correlation model in [wu2023multiple], Kosasih et al[kosasih2024finite] introduced the concept of a finite beam depth for spatial multiplexing in the distance domain. In the conventional far‐field regime, the beam depth tends to diverge to infinity—resulting in very high correlation between collinear users and hence precluding distance‐domain multiplexing. By contrast, the near‐field region can yield a finite beam depth, which is equivalent to low spatial correlation between two collinear users, thereby enabling spatial multiplexing. This finite beam depth has been characterized analytically for a variety of aperture geometries, including rectangular and circular apertures [kosasih2024finite], modular apertures [kosasih2024achieving, li2024multi], and sparse arrays [chen2024near, zhou2024sparse], highlighting the potential for distance‐domain spatial multiplexing by using various types of apertures.
@@
To obtain an closed-form expression of the distance-domain DoF, we consider an equivalent point-to-point MIMO (P2P-MIMO) channel between a base station as a continuous-aperture transmit array of arbitrary shape and a continuous-aperture arbitrarily long receive array that spans solely over the distance domain. The continuous array can be arbitrarily long, capturing scenarios where users span a wide range of distances. Moreover, we allow the base-station aperture to have arbitrary shape, reflecting practical design flexibility. We focus on the LoS component, which typically dominates in mmWave/THz bands due to significant NLoS attenuation [priebe2013stochastic]. Since multipath scattering typically increases channel rank by introducing virtual sources [tse2005fundamentals], our LoS analysis establishes a fundamental DoF lower bound characterizing performance in sparse-scattering environments.
I-A Related Works
@@
The DoF of near-field channels have been extensively investigated in point-to-point (P2P) MIMO settings; therefore, it is important to delineate how our problem relates to and differs from that literature. In general, most analyses consider geometries in which the receive (Rx) aperture spans primarily the angular domain of the transmit (Tx) aperture. Prior work has largely focused on configurations where one aperture covers the angular extent of the other, including linear arrays [xie2023performance, pu2014effects, torkildson2011indoor], rectangular arrays [hu2018beyond, dardari2020communicating, sun2021small, pizzo2022landau, do2023parabolic, gong2024holographic, wang2025analytical, guo2025impact], circular array [xu2017degrees] and more general apertures [yuan2021electromagnetic, shen2023electromagnetic, xu2023exploiting, miller2000communicating, pizzo2022nyquist, gustafsson2025degrees].
@@
In contrast to this extensive treatment of angular-domain DoF, the distance-domain DoF has received comparatively little attention. Some works do consider cases where the Rx aperture varies mainly along the distance domain of the Tx array [decarli2021communication, ruiz2023degrees, ding2022degrees, chen2025degrees]. In [decarli2021communication, ruiz2023degrees], the DoF between two rectangular apertures with mutually perpendicular orientations is analyzed. Although the Rx aperture in these studies spans the distance domain, it also extends over the angular domain. Hence, the DoF expression in [decarli2021communication, ruiz2023degrees] is a combination of the distance-domain and angular-domain DoF, which prevents us from understanding the true values of the distance-domain DoF. In [ding2022degrees, Sec. IV-B], two linear arrays with perpendicular orientations are considered; this setup can capture distance-domain DoF for collinear users when the base station is modeled as a linear array, but it does not generalize to arbitrary Tx aperture shapes. Moreover, [chen2025degrees] derives a closed-form DoF expression for a rectangular Tx aperture and two collinear users at fixed positions. While insightful, the restriction to two predetermined users does not address the broader question studied here: determining the distance-domain DoF, equivalently maximum number of collinear users that can be spatially multiplexed solely via distance separation. Finally, all the works in [decarli2021communication, ding2022degrees, ruiz2023degrees, chen2025degrees] assume Tx–Rx separations larger than the arrays dimensions, which does not reflect scenarios with collinear users distributed over a wide distance span.
@@ I-B3 Distinction from Angular-Domain DoF Results
The remainder of the paper is organized as follows. In Section II, we describe the system model, including the geometrical setup of Tx and Rx arrays, the near-field channel model. Section III presents the definition of distance-domain DoF and its implications in practice. Section IV presents the mathematical framework and derives a closed-form expression for the distance-domain DoF in the ideal case of Tx array and Rx array, with Rx array located in the broadside of Tx array  [^2] Parts of this paper including Section II, III-A and IV was presented in IEEE Globecom 2025 [son2025].. Section V extends the analysis to the non-broadside configuration of the Rx array. Section VI extends the analysis to the modular Tx array. Finally, Section VII concludes the paper.
@@ II-A Geometrical Setup
We consider a LoS channel between a base station (BS) equipped with a two-dimensional surface transmit (Tx) aperture and a receive (Rx) linear array that represents multiple collinear users. The coordinate origin $\mathbf{O}=[0,0,0]^{\mathsf{T}}$ is the intersection of the Tx-aperture plane and the Rx-array axis. The geometrical details are as follows.
@@ III-A Distance-domain DoF in Point-to-point MIMO Channel
III-B Implications of Distance-Domain DoF for MU–MIMO and Channel Estimation
@@ III-B1 Upper bound on MU–MIMO DoF (same angle, distance-separated)
$\mathbf{V}$ $\approx\mathbf{E}\,\mathbf{\Lambda}\,\mathbf{E}^{\mathsf{H}},$ (17)
$\mathbf{E}$ $=[\,\mathbf{e}_{1}\ \cdots\ \mathbf{e}_{N}\,],\ \mathbf{e}_{n}=[\,e_{n}(\mathbf{q}_{1}),\ldots,e_{n}(\mathbf{q}_{K})\,]^{\mathsf{T}},$ (18)
$\mathbf{\Lambda}$ $=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N}).$ (19)
@@
$\mathrm{rank}(\mathbf{E}\,\mathbf{\Lambda}\,\mathbf{E}^{\mathsf{H}})\leq\min\{\mathrm{rank}(\mathbf{E}),\,\mathrm{rank}(\mathbf{\Lambda})\}\leq N\leq N_{\varepsilon}.$ (20)
Therefore, the CAP DoF upper-bounds the DoF of any discrete MU–MIMO realization sampled along that angle. If there are more users than the available distance-domain DoF between $r_{\min},r_{\max}]$ $(K>N_{\varepsilon})$, the channel matrix $\mathbf{H}$ necessarily has small singular values, leading to ill-conditioning and degraded spatial multiplexing gain of MMSE/ZF-type precoders whose performance depends on the singular-value distribution.
@@ IV-A Formulation of Integral Operator with Convolution Kernel
We cannot apply Landau’s Theorem [landau1975szego] into the operator $\mathcal{V}$ since its kernel ${v}(r,r^{\prime})$ in (27) is not convolutional. This is because the amplitude ${1}/{rr^{\prime}}$ is not constant and $r$ can vary over a large distance. This is different from other P2P MIMO channel, where the element-wise distance between Tx and Rx arrays are relatively constant. To address this problem, a simple trick is to perform the following variable change:
$t=\frac{1}{r}\quad\text{and}\quad t^{\prime}=\frac{1}{r^{\prime}}.$ (29)
@@
$(\mathcal{G}\Phi)(t)=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}g(t,t^{\prime})\,\Phi(t^{\prime})\,dt^{\prime},\quad t\in\Big[r^{-1}_{\max},r^{-1}_{\min}\Big],$ (30)
with $g(t,t^{\prime})$ being the convolution kernel:
$g(t,t^{\prime})=g(t-t^{\prime})$ $=\iint_{\mathcal{P}}\exp\!\Big(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}(t-t^{\prime})\Big)d\mathbf{p},$
$\quad\quad\quad\quad\quad\quad\forall\,t,t^{\prime}\in\Big[r^{-1}_{\max},r^{-1}_{\min}\Big].$ (31)
@@
Lemma 1 indicates that the number of dominant eigenvalues of $\mathcal{V}$} can be deduced by analyzing $\mathcal{G}$. Moreover, because $g(t,t^{\prime})$ is a Hermitian convolution kernel, the number of dominant eigenvalues of $\mathcal{G}$ can be inferred from Landau’s Theorem (or Szego’s Theorem) [landau1975szego]:
$\text{DoF}=\text{Bandwidth}\big(\hat{g}(f)\big)\times T,$ (32)
where $\hat{g}(f)=\int_{-\infty}^{\infty}g(t)e^{-j2\pi ft}dt$ is the Fourier transform of $g(t)$, $T={r^{-1}_{min}}-{r^{-1}_{max}}$ represents the length of the interval in the inverse-distance domain. Intuitively, the operator $\mathcal{G}$ in (30) with convolution kernel $g(t)$ behaves like a linear time-invariant system in which an input signal of duration $T$ is transmitted through a channel with frequency response $g(t)$. Thus, the DoF of this system corresponds to the number of data symbols that can be transmitted in this channel in duration $T$. In our setting, the time domain is replaced by the inverse-distance domain, and the frequency domain becomes the inverse-distance frequency domain.
For simplicity, we rescale $\hat{g}(f)$ by $T$ (i.e., setting $\xi=Tf$). Hence, the DoF is simply equal to the bandwidth of $\tilde{g}(\xi)$ where:
$\tilde{g}(\xi)=\hat{g}\!\left(\frac{\xi}{T}\right)=\int_{-\infty}^{\infty}g(t)e^{-j2\pi\xi\frac{t}{T}}dt.$ (33)

### v4

@@ Abstract
Extremely large aperture array operating in the near‐field regime unlock additional spatial resources, which can be exploited to simultaneously serve multiple users even when they share the same angular direction. This work investigate the distance-domain degrees of freedom (DoF), which are the DoF when the user only varies its distance to the base station and not the angle. To obtain the distance-domain DoF, we investigate a line-of-sight (LoS) channel between a base station (source) and observation region representing users. The base station is modeled as a large two-dimensional transmit (Tx) array with an arbitrary shape. The observation region is modeled as an arbitrarily long linear receive (Rx) array, where elements are collinearly aligned but located at varying distances from the Tx array. We assume that both the Tx and Rx array has continuous aperture with an infinite number of elements and infinitesimal spacing, which establishes an upper bound for the distance-domain DoF in the case of finite number of elements. First, we assume an ideal case where the Tx array is a single piece and the Rx array is on the broad side of the Tx array. By reformulating the channel as an integral operator with a Hermitian convolution kernel, we derive a closed-form expression for the distance-domain DoF via the Fourier transform. Our analysis shows that the distance-domain DoF is predominantly determined by the extreme boundaries of both the Tx and Rx arrays rather than its detailed interior structure. We further extend the framework to non-broadside configurations by employing a projection method that converts the problem to an equivalent broadside case. Finally, we extend our analytical framework to the modular array, which shows the distance-domain DoF gain over the single-piece array given the constraint of the physical length of the array.
@@ I Introduction
Unlike far-field scenarios, where beamforming is mainly used to steer energy along specific angular directions, the near-field enables beam focusing that leverages both angular and distance domains. This capability permits spatial multiplexing of users that share the same angular direction if they are located at different distances. Such an approach could be particularly useful in ultra-dense scenarios where many users are aligned in the same direction, a situation that challenges traditional far-field beamforming methods. Several studies have investigated spatial multiplexing based on the distance domain; for instance, [zhang2022beam] and [li2023near] demonstrated through simulations that two collinear users can be spatially multiplexed in a LoS near-field channel. However, these works did not delve into a theoretical explanation of channel characteristics for the effectiveness of such multiplexing. To address this gap, [wu2023multiple] and [wu2023enabling] investigated the channel correlation between collinear users in the near-field region. Once the correlation falls below a certain threshold, spatial multiplexing becomes feasible. Based on the closed‐form channel‐correlation model in [wu2023multiple], Kosasih et al[kosasih2024finite] introduced the concept of a finite beam depth for spatial multiplexing in the distance domain. In the conventional far‐field regime, the beam depth tends to diverge to infinity—resulting in very high correlation between collinear users and hence precluding distance‐domain multiplexing. By contrast, the near‐field region can yield a finite beam depth, which is equivalent to low spatial correlation between two collinear users, thereby enabling spatial multiplexing. This finite beam depth has been characterized analytically for a variety of aperture geometries, including rectangular and circular apertures  [kosasih2024finite], modular apertures [kosasih2024achieving, li2024multi], and sparse arrays [chen2024near, zhou2024sparse]. These studies highlight the strong potential for distance-domain spatial multiplexing across diverse antenna architectures.
@@
To obtain a closed-form expression of the distance-domain DoF, we consider an equivalent point-to-point MIMO (P2P-MIMO) channel between a base station as a continuous-aperture transmit array of arbitrary shape and a continuous-aperture arbitrarily long receive array that spans solely over the distance domain. The continuous array can be arbitrarily long, capturing scenarios where users span a wide range of distances. Moreover, we allow the base-station aperture to have arbitrary shape, reflecting practical design flexibility. We focus on the LoS component, which typically dominates in mmWave/THz bands due to significant NLoS attenuation [priebe2013stochastic]. Since multipath scattering typically increases channel rank by introducing virtual sources [tse2005fundamentals], our LoS analysis establishes a fundamental DoF lower bound characterizing performance in sparse-scattering environments.
I-A Related Work
@@
The DoF of near-field channels have been extensively investigated in point-to-point (P2P) MIMO settings; therefore, it is important to delineate how our problem relates to and differs from that literature. In general, most analyses consider geometries in which the receive (Rx) aperture spans primarily the angular domain of the transmit (Tx) aperture. Prior work has largely focused on configurations where one aperture covers the angular extent of the other, including linear arrays [xie2023performance, pu2014effects, torkildson2011indoor, kanatas2025deterministic], rectangular arrays [hu2018beyond, dardari2020communicating, sun2021small, pizzo2022landau, do2023parabolic, gong2024holographic, wang2025analytical, guo2025impact, hamza2025spatial, hussain2026analyzing], circular array [xu2017degrees] and more general apertures [yuan2021electromagnetic, shen2023electromagnetic, xu2023exploiting, miller2000communicating, pizzo2022nyquist, gustafsson2025degrees]. Mathematically, spatial DoF evaluation in near-field configurations heavily relies on integral operator kernels. The literature primarily utilizes two kernel-based approaches:
  1. 1.
    Estimating the effective DoF via the trace of the spatial correlation kernel—typically an energy ratio between its squared trace and Frobenius norm [xie2023performance, wang2025analytical, yuan2021electromagnetic, shen2023electromagnetic, gustafsson2025degrees, hamza2025spatial, hussain2026analyzing].
  2. 2.
    Exploiting the band-limited characteristics of the communication kernel via Landau’s eigenvalue theorem [pizzo2022landau, do2023parabolic] or Prolate Spheroidal Wave Functions (PSWFs) [pu2014effects, torkildson2011indoor, guo2025impact].
While these studies successfully apply kernel methods for standard user distributions, isolating the pure distance-domain DoF for collinear users requires a fundamentally different kernel transformation, which is the focus of this work.
@@
Compared to this extensive treatment of angular-domain DoF, the distance-domain DoF has received comparatively little attention. Some works do consider cases where the Rx aperture varies mainly along the distance domain of the Tx array [decarli2021communication, ruiz2023degrees, ding2022degrees, chen2025degrees]. In [decarli2021communication, ruiz2023degrees], the DoF between two rectangular apertures with mutually perpendicular orientations is analyzed. Although the Rx aperture in these studies spans the distance domain, it also extends over the angular domain. Hence, the DoF expression in [decarli2021communication, ruiz2023degrees] is a combination of the distance-domain and angular-domain DoF, which prevents us from understanding the true values of the distance-domain DoF. In [ding2022degrees, Sec. IV-B], two linear arrays with perpendicular orientations are considered; this setup can capture distance-domain DoF for collinear users when the base station is modeled as a linear array, but it does not generalize to arbitrary Tx aperture shapes. Moreover, [chen2025degrees] derives a closed-form DoF expression for a rectangular Tx aperture and two collinear users at fixed positions. While insightful, the restriction to two predetermined users does not address the broader question studied here: determining the distance-domain DoF, equivalently maximum number of collinear users that can be spatially multiplexed solely via distance separation. Finally, all the works in [decarli2021communication, ding2022degrees, ruiz2023degrees, chen2025degrees] assume Tx–Rx separations larger than the arrays dimensions, which does not reflect scenarios with collinear users distributed over a wide distance span.
@@ I-B3 Distinction from Angular-Domain DoF Results
The remainder of the paper is organized as follows. In Section II, we describe the system model, including the geometrical setup of Tx and Rx arrays, the near-field channel model. Section III presents the definition of distance-domain DoF and its implications in practice. Section IV presents the mathematical framework and derives a closed-form expression for the distance-domain DoF in the ideal case of Tx array and Rx array, with Rx array located in the broadside of Tx array  [^2] Parts of this paper including Section II, III-A and IV was presented in IEEE Globecom 2025 [duong2025spatial].. Section V extends the analysis to the non-broadside configuration of the Rx array. Section VI extends the analysis to the modular Tx array. Finally, Section VII concludes the paper.
@@ II-A Geometrical Setup
To rigorously isolate the distance-domain DoF from the angular-domain DoF, our LoS MIMO channel assumes all receive (Rx) elements align along a single angle, differing only in radial distance. This mathematically reduces the observation region to a 1D linear array, because a 2D Rx region would inherently entangle the distance- and angular-domain DoF. Conversely, we assume an arbitrary 2D shape for the transmit (Tx) array to maximize the generality of our analytical framework across diverse base station architectures. The detailed system parameters are defined as follows.
@@ III-A Distance-domain DoF in Point-to-point MIMO Channel
In practical systems, the choice of $\epsilon$ is determined by the environmental noise floor, ensuring that the spatial sub-channels associated with the counted eigenvalues have sufficient strength for reliable communication. Additionally, the effective DoF calculation is notably stable with respect to the choice of $\epsilon$. Because the near-field eigenvalues experience a exponential decay following the cutoff index, adjusting the threshold from $\epsilon_{1}$ to $\epsilon_{2}$ results in merely a marginal logarithmic shift, scaling as $\mathcal{O}(\log(\epsilon_{2}/\epsilon_{1}))$.
III-B Implications of Distance-Domain DoF for MU-MIMO and Channel Estimation
@@ III-B1 Upper bound on MU–MIMO DoF (same angle, distance-separated)
$\mathbf{V}$ $\approx\mathbf{E}\mathbf{\Lambda}\mathbf{E}^{\text{H}},$ (17)
$\mathbf{E}$ $=[\,\mathbf{e}_{1}\ \cdots\ \mathbf{e}_{N}\,],\ \mathbf{e}_{n}=[\,e_{n}(\mathbf{q}_{1}),\dots,e_{n}(\mathbf{q}_{K})\,]^{\text{T}},$ (18)
$\mathbf{\Lambda}$ $=\text{diag}(\lambda_{1},\dots,\lambda_{N}).$ (19)
@@
$\text{rank}(\mathbf{E}\mathbf{\Lambda}\mathbf{E}^{\text{H}})\leq\min\{\text{rank}(\mathbf{E}),\text{rank}(\mathbf{\Lambda})\}\leq N\leq N_{\varepsilon}.$ (20)
Therefore, the CAP DoF upper-bounds the DoF of any discrete MU–MIMO realization sampled along that angle. If there are more users than the available distance-domain DoF between $r_{\min},r_{\max}]$ $(K>N_{\varepsilon})$, the channel matrix $\mathbf{H}$ necessarily has small singular values, leading to ill-conditioning and degraded spatial multiplexing gain of MMSE/ZF-type precoders whose performance depends on the singular-value distribution. Consequently, this rank deficiency imposes a strict physical ceiling on the achievable multi-user sum-capacity; because inter-user interference cannot be effectively mitigated in the spatial domain when $K>N_{\varepsilon}$, the system capacity will inevitably saturate or degrade.
@@ IV-A Formulation of Integral Operator with Convolution Kernel
Note that its kernel ${v}(r,r^{\prime})$ in (27) is not Hermitian kernel, since the amplitude ${1}/{rr^{\prime}}$ is not constant and $r$ can vary over a large distance, and the phase difference is a function of $1/r-1/r^{\prime}$ rather than the shift-invariant difference $1/r-1/r^{\prime}$. This is different from other P2P MIMO channel, where the element-wise distance between Tx and Rx arrays are relatively constant. To facilitate the DoF analysis via Landau’s Theorem [landau1975szego], we perform the following variable change:
$\zeta=\frac{1}{r}\quad\text{and}\quad\zeta^{\prime}=\frac{1}{r^{\prime}}.$ (29)
@@
$(\mathcal{G}\Phi)(\zeta)=\int_{r^{-1}_{\max}}^{r^{-1}_{\min}}g(\zeta,\zeta^{\prime})\,\Phi(\zeta^{\prime})\,dt^{\prime},\quad\zeta\in\Big[r^{-1}_{\max},r^{-1}_{\min}\Big],$ (30)
with $g(\zeta,\zeta^{\prime})$ being the Hermitian convolutional kernel:
$g(\zeta,\zeta^{\prime})=g(\zeta-\zeta^{\prime})$ $=\iint_{\mathcal{P}}\exp\!\Big(-\frac{j2\pi}{\lambda}\frac{\|\mathbf{p}\|^{2}}{2}(\zeta-\zeta^{\prime})\Big)d\mathbf{p},$
$\quad\quad\quad\quad\quad\quad\forall\,\zeta,\zeta^{\prime}\in\Big[r^{-1}_{\max},r^{-1}_{\min}\Big].$ (31)
Crucially, because the spatial integration over $\mathbf{p}\in\mathcal{P}$ is decoupled from the distance variables, the kernel $g(\zeta,\zeta^{\prime})$ inherently depends only on the linear difference $(\zeta-\zeta^{\prime})$. Consequently, it remains a valid Hermitian convolution kernel regardless of the specific shape of the arbitrary bounded aperture $\mathcal{P}$.
@@
Lemma 1 indicates that the number of dominant eigenvalues of $\mathcal{V}$} can be deduced by analyzing $\mathcal{G}$. Moreover, because $g(\zeta,\zeta^{\prime})$ is a Hermitian convolution kernel, the number of dominant eigenvalues of $\mathcal{G}$ can be inferred from Landau’s Theorem (or Szego’s Theorem) [landau1975szego]:
$\text{DoF}=\text{Bandwidth}\big(\hat{g}(f)\big)\times L_{\zeta},$ (32)
where $\hat{g}(f)=\int_{-\infty}^{\infty}g(\zeta)e^{-j2\pi f\zeta}dt$ is the Fourier transform of $g(\zeta)$, $L_{\zeta}={r^{-1}_{min}}-{r^{-1}_{max}}$ represents the length of the interval in the inverse-distance domain. Intuitively, the operator $\mathcal{G}$ in (30) with convolution kernel $g(\zeta)$ behaves like a linear time-invariant system in which an input signal of duration $L_{\zeta}$ is transmitted through a channel with frequency response $g(\zeta)$. Thus, the DoF of this system corresponds to the number of data symbols that can be transmitted in this channel in duration $L_{\zeta}$. In our setting, the time domain is replaced by the inverse-distance domain, and the frequency domain becomes the inverse-distance frequency domain.
For simplicity, we rescale $\hat{g}(f)$ by $L_{\zeta}$ (i.e., setting $\xi=L_{\zeta}f$). Hence, the DoF is simply equal to the bandwidth of $\tilde{g}(\xi)$ where:
$\tilde{g}(\xi)=\hat{g}\!\left(\frac{\xi}{L_{\zeta}}\right)=\int_{-\infty}^{\infty}g(\zeta)e^{-j2\pi\xi\frac{\zeta}{L_{\zeta}}}d\zeta.$ (33)